QUESTION 2 Earthquakes produce several types of shock waves, two of which are called S and P waves. P waves travel at about 6.5 km/s, and S waves travel at about 3.5 km/s. The delay in arrival times between P waves and S waves provides from the measurement station. If a station records a delay of 33 s, how far from the station was the earthquake (ín kilometers? Use two significant figures in your answer. information about how far away the earthquake epicenter is QUESTION 3 An important piece of landing equipment must be thrown to a ship, which is moving at 45 cm/s towards the shore, before the ship can dock. This equipment is thrown at 15 m/s at 60° above the horizontal from the top of a tower at the edge of the water, 8.75 m above the ships deck as shown. For this equipment to land at the front of the ship, what distance D from the dock, in meters, should the ship be when the equipment is thrown? Ignore air resistance. QUESTION 4 A subway train starts from rest at a station and accelerates along a straight track at 1.60 m/sf for 14 S. It then runs at constant speed for 70.0 s, and then slows down at a rate of 3.50 m/s2 until it stops at the next station. Find the total distance, in meters, traveled between the stations.
Answer #1

Question 2)

Speed of P waves UP = 6.5 km/s

Speed of S waves us= 3.5km/s

Distance of station from earthquake is d

Time delay between P and S waves is ts-tp-33s

v_p=rac{d}{t_p}Rightarrow t_p=rac{d}{v_p} , and   v_s=rac{d}{t_s}Rightarrow t_s=rac{d}{v_s}

ts -tp ひ

15-3) 3.5 6.5

3 33 d d = 250.25 km 3,5 * 6.5

Distance of station from earthquake is d250,25 km

--------------

Question 3)

The equipment is at height 8.75,m,

When the equipment is thrown on to ship, the ship is at a distance of D from the dock.

D= Distance traveled by equipment towards the ship (in horizontal direction) + distance traveled by ship towards the dock.

Time taken by ship to move a distance x_2 is t. In the same time the equipment moves through a distance x_1.

Vertical distance traveled by equipment in time t is  y_1=-8.75,m

Using y_1=u_yt-rac{1}{2}gt^2

-8.75=15sin60degree t-rac{1}{2}(9.81)t^2

-4.905t^2+13,t+8.75=0

t 3.207s

Horizontal distance traveled by equipment in time t is x_1=u_x t=15cos60degree *3.207=24.05,m

Distance covered by ship during this time is x_2=v_s t=0.45*3.207=1.44,m

D=x_1+x_2=24.05+1.44=25.5,m

---------

Question 4)

Train starts with an initial speed  v_1=0,m/s and accelerates for t_1=14,s with an acceleration a_1=1.60,m/s^2

Distance covered during this time S_1=v_1t_1+rac{1}{2}a_1t_1^2=0+rac{1}{2}*(1.60)(14)^2=156.8,m

v_2=v_1+a_1t_1=0+1.60*14=22.4,m/s

For the next t_2=70,s the train travels with constant velocity v_2

Distance covered during the time t_2 is  S_s=v_2t_2+rac{1}{2}a_2t_2^2=22.4*70+0=1568,m

Then train slows down to v_3=0,m/swith an acceleration of a_3=-3.50,m/s^2 in time t_3

Using v_3^2-v_2^2=2a_3S_3

Distance covered during this interval is  S_3=rac{v_3^2-v_2^2}{2a_3}=rac{0^2-22.4^2}{2*(-3.50)}=71.68,m

Total distance covered between the stations is  is S-S1S2+S3-1796.5m

------------------